OpenStudy (anonymous):

PLEASE HELP can someone help me solve this quadratic equation? x2 = −2x + 24

OpenStudy (agreene):

What aren't you understanding?

OpenStudy (jim766):

Add 2x and sub 24 on both sides, then factor

OpenStudy (anonymous):

i just need help solving it like the equation to solve it

OpenStudy (agreene):

\[x^2+2x=24\] \[x^2+2x+1=25\] \[(x+1)^2=25\] \[x=\{4,-6\}\]

OpenStudy (anonymous):

oh thank you!!!!!!!!!!!!!